# Dynamic modes of active Potts models with factorizable numbers of states

### v1

@@ III.1.1 Modes for $J_{k,[k+3]}=J_{k,[k+2]}$
For second-order transitions, the scaling relations of the order parameter (here, $R_{3}$) are obtained in the vicinity of the transition point $J_{\mathrm{c}}$: $R_{3}L^{\beta/\nu}=\tilde{R}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}}]$, $\chi_{3}L^{-\gamma/\nu}=\tilde{\chi}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}})]$, and $U_{3}=\tilde{U}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}})]$, where $\tilde{R}_{3}(x)$, $\tilde{\chi}_{3}(x)$, and $\tilde{U}_{3}(x)$ are the scaling functions and $\nu$, $\beta$, and $\gamma$ are the critical exponents. Figure 6 shows the normalized $U_{3}$, $\chi_{3}$, and $R_{3}$ as functions of $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}$ for $h=0$ and $1$. We fitted each curve with a quartic function and calculated the scaling exponents by least-squares fits to these functions for $L\leq 90$ in the vicinity of the transition point. We obtained the following values: $J_{\mathrm{c}}=1.725\pm 0.002$, $1/\nu=0.95\pm 0.02$, $\gamma/\nu=1.742\pm 0.002$, and $\beta/\nu=0.1233\pm 0.0004$ for $h=0$; and $J_{\mathrm{c}}=1.214\pm 0.001$, $1/\nu=1.12\pm 0.02$, $\gamma/\nu=1.732\pm 0.007$, and $\beta/\nu=0.132\pm 0.002$ for $h=1$. The exponent $\nu$ is close to unity and is smaller for $h=1$ than for $h=0$ (equilibrium). Therefore, the exponents for the transition from M6 to the wave mode (W3) are modified from the equilibrium values between the mixed phases. Note that, in Fig. 6, for $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}\gtrsim 10$, the curves do not converge well either for $h=0$ or $1$. This deviation may be caused by the crossover effects [21] from other transitions (e.g., the W6–W3 transition).
Figure 10: Transition between HC3 and M3 modes at $J_{k,[k+3]}=J_{k,[k+2]}$ for (a–c) $h=0$ and (d–f) $h=1$. Binder cumulant $U_{1}$, susceptibility $\chi_{1}$, and order parameter $R_{1}$ on the single-phase occupancy are shown in (a,d), (b,e), and (c,f), respectively. The transition points are $J_{\mathrm{c}}=0.992$ and $0.977$ for $h=0$ and $1$, respectively.
@@ III.1.2 Modes for $J_{k,[k+3]}=0$
Next, we investigate the $J_{k,[k+2]}$ dependence while keeping $J_{k,[k+3]}=0$ (Figs. 710). Although the dynamics for $J_{k,[k+2]}\leq 0$ are similar to those for $J_{k,[k+3]}=J_{k,[k+2]}$, two new modes (HC3 and M3) are found for $J_{k,[k+2]}>0$, instead of the W3 and M6 modes.
At $J_{k,[k+2]}=J_{k,k}=2$, three odd-numbered states ($s=1,3,5$) can be randomly mixed without lose of contact energy in the absence of the even-numbered states; similarly for the even-numbered states. By contrast, mixing of the six states is inhibited by the repulsion between the diagonal pairs ($s=k$ and $[k+3]$) because $J_{k,[k+3]}=0$. Thus, the mixed phases of the three states (M3) are stable for $J_{k,[k+2]}\simeq 2$, for which the boundaries of the clusters of each state are not smooth and moves diffusively rather than ballistically (see Figs. 7(b), (c), and 8(b), and Movie S3). No switching between the odd- and even-numbered states was observed during the available simulation periods $\sim 10^{9}$. This phase also exists at $h=0$.
@@
The transition between the HC3 and M3 modes is continuous. As $J_{k,[k+2]}$ increases in the HC3 range, the densities of the dominant phase and the other two phases of the three major states decreases and increase, respectively (see Fig. 8). Additionally, the contact probability $p_{\mathrm{contact}}$ between these three states increases (see Fig. 9(b)). Hence, the three densities converge as the system approaches the M3 mode. This change can be characterized by the order parameter $R_{1}$. We calculated the scaling exponents using the method described in Sec. III.1.1. Figure 10 shows the normalized $U_{1}$, $\chi_{1}$, and $R_{1}$ as functions of $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}$ for $h=0$ and $1$. We obtained the following values: $J_{\mathrm{c}}=0.992\pm 0.001$, $1/\nu=1.24\pm 0.05$, $\gamma/\nu=1.75\pm 0.02$, and $\beta/\nu=0.137\pm 0.002$ for $h=0$; and $J_{\mathrm{c}}=0.977\pm 0.001$, $1/\nu=1.19\pm 0.04$, $\gamma/\nu=1.761\pm 0.008$, and $\beta/\nu=0.126\pm 0.005$ for $h=1$. These values are close to the $\nu=5/6$, $\gamma=13/9$, and $\beta=1/9$ obtained for the three-state standard Potts model [16]. The differences in the exponents for $h=0$ and $1$ are in the ranges of statistical errors, thus, the scaling exponents are not or little modified from the equilibrium values. This outcome may reflect the static nature of the two modes. That is, in most periods of the HC3 mode, a single-state dominant phase exists, similar to that obtained at equilibrium for $h=0$. The M3 mode is almost identical to the mixed phase at equilibrium. To obtain a distinct change in the scaling exponents, dynamic modes, such as wave propagation in W3, may be required.
@@
Figure 11: Six-state Potts model at $J_{k,[k+2]}=0$ and $L=128$. (a) Snapshot of spiral waves of three types of two-state mixed domains (M2W3) at $h=1.3$ and $J_{k,[k+3]}=2$. (b) Snapshot of homogeneous cycling of mixed phases (M2HC3) at $h=1.3$ and $J_{k,[k+3]}=2.5$. Two diagonal states ($s=2$ and $5$) are mixed. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, green down-pointing triangles, and light-blue squares represent HC6, M2W3, M2HC3, W6, WI, and W3, respectively.
Figure 12: Dependence on $J_{k,[k+3]}$ in the six-state Potts model at $J_{k,[k+2]}=0$ and $h=1.3$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, and $k$$[k+3]$ pairs. (c) Binder cumulant $U_{3}$ for three-fold symmetry and cluster size ratio $d_{\mathrm{cl}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=256$.
Figure 13: Time development of number densities of states in the six-state Potts model at $J_{k,[k+2]}=0$ and $L=128$. (a) Temporal coexistence of M2W3 and M2HC3 at $h=1.3$ and $J_{k,[k+3]}=2.4$. (b) Temporal coexistence of W3 and HC6 at $h=0.7$ and $J_{k,[k+23]}=-0.6$.
@@ III.1.3 Modes for $J_{k,[k+2]}=0$
Next, we investigate the $J_{k,[k+3]}$ dependence while keeping $J_{k,[k+2]}=0$ (Figs. 1113). For $J_{k,[k+3]}<0$, the W3 mode emerges in addition to the HC6, WI, and W6 modes, since the contact between the $k$ and $[k+2]$ states becomes more stable than that between the $k$ and $[k+3]$ states (see Figs. 11(c) and 12). The transition between the W3 and HC6 modes occurs via the temporal coexistence of the two modes, as shown in Fig. 13(b), similar to that occurring for the transition from HC3 to W3 in the three-state Potts model [34, 35]. The ratio of the W3 modes monotonously increases with decreasing $J_{k,[k+3]}$, and we determine the modes from the ratio of the phase periods.
[…]
In the M2W3 mode, the three types of mixed domains form spiral waves (see Fig. 11(a) and Movie S4). To characterize this mode, we calculated the $p_{\mathrm{contact}}$ values and cluster size ratio $d_{\mathrm{cl}}$. Clusters of single or diagonal two states are considered. Note that, in the latter case, when neighboring sites have the same state or diagonal state ($s_{i}=k$ and $s_{j}=k$ or $[k+3]$), they belong to the same cluster. With increasing $J_{k,[k+3]}$, the ratio $d_{\mathrm{cl}}=n_{\mathrm{c1}}/n_{\mathrm{c2}}$ of the mean sizes for these two clusters ($n_{\mathrm{c1}}$ and $n_{\mathrm{c2}}$ for the single and two diagonal states, respectively) decreases and $p_{\mathrm{contact}}$ of the diagonal pairs ($s=k$ and $[k+3]$) increases (see Fig. 12(b) and (c)). We consider that the W6 mode changes to the M2W3 mode when the the diagonal-pair $p_{\mathrm{contact}}$ exceeds the other probabilities.
Figure 14: Dependence on $J_{k,[k+3]}-J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+2]}+J_{k,[k+3]}=1$, $h=1$, and $L=128$. (a) Order parameters $R_{1}$ and $R_{3}$. The bidirectional arrows at the top represent the ranges of dynamic modes, including the coexistence states. (b) Susceptibilities $\chi_{1}$ and $\chi_{3}$. The conditions of $J_{k,[k+3]}-J_{k,[k+2]}=-1$, 00, and $1$ correspond to the points in the phase diagram of Figs. 7, 1, and 11, respectively. (c) Time development of number densities of states at $J_{k,[k+3]}-J_{k,[k+2]}=-0.4$. The HC3 and W3 modes temporally coexist.
Figure 15: Probabilities $p_{\mathrm{phase}}$ of one-phase dominant and multi-phase coexistence states as functions of $h$ at $J_{k,k^{\prime}}=2\delta_{k,k^{\prime}}$ and $L=128$ for the $q$-state Potts models with (a) $q=8$, (b) $q=5$, and (c) $q=4$. The plot in (c) is reproduced from Ref. 36. Licensed under CC BY.
@@ III.2 Five-State Potts Model
[…]
The dynamic phase diagram of the five-state Potts model is shown in Fig. 16. Since five is a prime number, no factorized-symmetry modes exist. At $J_{k,[k+2]}\simeq J_{k,k}=2$, a five-state mixed phase (M5) exists, as for the six-state Potts model. To characterize the transition between the M5 and W5 modes, we calculated $p_{\mathrm{phase}}$, $p_{\mathrm{contact}}$, and the mean cluster size $n_{\mathrm{c1}}$, as shown in Fig. 17; At $J_{k,[k+2]}\simeq 0.6$, a stepwise change is observed for $p_{\mathrm{phase}}$ of five states. Additionally, a minimum is observed for $p_{\mathrm{contact}}$ of $k$$[k+1]$ pairs and a sharp decrease in $n_{\mathrm{c1}}$ is initiated. For $J_{k,[k+2]}\lesssim 0.6$, $n_{\mathrm{c1}}$ is dependent on the system size; however, no size dependence is detected for $J_{k,[k+2]}\gtrsim 0.6$. In contrast, no size dependence is obtained for $p_{\mathrm{contact}}$ in either range (compare the curves for $L=128$ and $256$ in Fig. 17(b) and (c)). Thus, we consider $J_{k,[k+2]}\simeq 0.6$ is the transition point for $h=1$. The minimum of $p_{\mathrm{contact}}$ is used to determine the transition points for other values of $h$, since it is clearer than the other two features.
@@
Figure 17: Dependence on $J_{k,[k+2]}$ in the five-state Potts model at $h=1$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$ and $k$$[k+2]$ pairs. (c) Mean cluster size $n_{\mathrm{c1}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=256$.

### v2

@@ III.1.1 Modes for $J_{k,[k+3]}=J_{k,[k+2]}$
For second-order transitions, the scaling relations of the order parameter (here, $R_{3}$) are obtained in the vicinity of the transition point $J_{\mathrm{c}}$: $R_{3}L^{\beta/\nu}=\tilde{R}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}}]$, $\chi_{3}L^{-\gamma/\nu}=\tilde{\chi}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}})]$, and $U_{3}=\tilde{U}_{3}[L^{1/\nu}(J_{k,[k+2]}-J_{\mathrm{c}})]$, where $\tilde{R}_{3}(x)$, $\tilde{\chi}_{3}(x)$, and $\tilde{U}_{3}(x)$ are the scaling functions and $\nu$, $\beta$, and $\gamma$ are the critical exponents. Figure 6 shows the normalized $U_{3}$, $\chi_{3}$, and $R_{3}$ as functions of $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}$ for $h=0$ and $1$. We fitted each curve with a quartic function and calculated the scaling exponents by least-squares fits to these functions for $L\leq 90$ in the vicinity of the transition point.
We obtained the exponents shown in Table 1 with $J_{\mathrm{c}}=1.725\pm 0.002$ and $1.214\pm 0.001$ for $h=0$ and $1$, respectively. The exponent $\nu$ is close to unity and is smaller for $h=1$ than for $h=0$ (equilibrium). Therefore, the exponents for the transition from M6 to the wave mode (W3) are modified from the equilibrium values between the mixed phases (M6 and M3). Note that, in Fig. 6, for $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}\gtrsim 10$, the curves do not converge well for either $h=0$ or $1$. This deviation may be caused by the crossover effects [21] from other transitions (e.g., the W6–W3 transition).
Table 1: Scaling exponents for the transitions between W3(M3) and M6 (Pot6a) and between HC3(H) and M3 (Pot6b) for $h=0$ and $1$.
$1/\nu$ $\gamma/\nu$ $\beta/\nu$
Pot6a
$h=0$ $0.95\pm 0.02$ $1.742\pm 0.002$ $0.1233\pm 0.0004$
$h=1$ $1.12\pm 0.02$ $1.732\pm 0.007$ $0.132\pm 0.002$
Pot6b
$h=0$ $1.24\pm 0.05$ $1.75\pm 0.02$ $0.137\pm 0.002$
$h=1$ $1.19\pm 0.04$ $1.761\pm 0.008$ $0.126\pm 0.005$
Figure 10: Transition between HC3 (H) and M3 modes at $J_{k,[k+3]}=0$ for (a–c) $h=0$ and (d–f) $h=1$ (Pot6b). Binder cumulant $U_{1}$, susceptibility $\chi_{1}$, and order parameter $R_{1}$ on the single-phase occupancy are shown in (a,d), (b,e), and (c,f), respectively. The transition points are $J_{\mathrm{c}}=0.992$ and $0.977$ for $h=0$ and $1$, respectively.
@@ III.1.2 Modes for $J_{k,[k+3]}=0$
Next, we investigate the $J_{k,[k+2]}$ dependence while keeping $J_{k,[k+3]}=0$ (Figs. 710). We refer to this condition as Pot6b. Although the dynamics for $J_{k,[k+2]}\leq 0$ are similar to those for $J_{k,[k+3]}=J_{k,[k+2]}$, two modes (HC3 and M3) are found for $J_{k,[k+2]}>0$, instead of the W3 and M6 modes.
At $J_{k,[k+2]}=J_{k,k}=2$, three odd-numbered states ($s=1,3,5$) can be randomly mixed without loss of contact energy in the absence of the even-numbered states; similarly for the even-numbered states. By contrast, mixing of the six states is inhibited by the repulsion between the diagonal pairs ($s=k$ and $[k+3]$) because $J_{k,[k+3]}=0$. Thus, the mixed phases of the three states (M3) are stable for $J_{k,[k+2]}\simeq 2$, for which the boundaries of the clusters of each state are not smooth and move diffusively rather than ballistically (see Figs. 7(b), (c), and 8(b), and Movie S3). No switching between the odd- and even-numbered states was observed during the available simulation periods $\sim 10^{9}$. This phase also exists at $h=0$.
@@
The transition between the HC3 and M3 modes is continuous. As $J_{k,[k+2]}$ increases in the HC3 range, the densities of the dominant phase and the other two phases of the three major states decreases and increase, respectively (see Fig. 8). Additionally, the contact probability $p_{\mathrm{contact}}$ between these three states increases (see Fig. 9(b)). Hence, the three densities converge as the system approaches the M3 mode. This change can be characterized by the order parameter $R_{1}$. At $h=0$ (equilibrium), it becomes the transition between the M3 and one-state dominant phases (called H). We calculated the scaling exponents using the method described in Sec. III.1.1. Figure 10 shows the normalized $U_{1}$, $\chi_{1}$, and $R_{1}$ as functions of $(J_{k,[k+2]}-J_{\mathrm{c}})L^{1/\nu}$ for $h=0$ and $1$. We obtained the exponents as shown in Table 1 with $J_{\mathrm{c}}=0.992\pm 0.001$ and $0.977\pm 0.001$ for $h=0$ and $1$, respectively. These values are close to the $\nu=5/6$, $\gamma=13/9$, and $\beta=1/9$ obtained for the three-state standard Potts model [16]. The differences in the exponents for $h=0$ and $1$ are in the ranges of statistical errors, thus, the scaling exponents are not or little modified from the equilibrium values. This outcome may reflect the static nature of the two modes. That is, in most periods of the HC3 mode, a single-state dominant phase exists, similar to that obtained at equilibrium for $h=0$. The M3 mode is almost identical to the mixed phase at equilibrium. To obtain a distinct change in the scaling exponents, dynamic modes, such as wave propagation in W3, may be required.
@@
Figure 11: Six-state Potts model at $J_{k,[k+2]}=0$ and $L=128$ (Pot6c). (a) Snapshot of spiral waves of three types of two-state mixed domains (M2W3) at $h=1.3$ and $J_{k,[k+3]}=2$. (b) Snapshot of homogeneous cycling of mixed phases (M2HC3) at $h=1.3$ and $J_{k,[k+3]}=2.5$. Two diagonal states ($s=2$ and $5$) are mixed. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, green down-pointing triangles, and light blue squares represent HC6, M2W3, M2HC3, W6, WI, and W3, respectively.
Figure 12: Dependence on $J_{k,[k+3]}$ in the six-state Potts model at $J_{k,[k+2]}=0$ and $h=1.3$ (Pot6c). (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for kk[k+1][k+1], kk[k+2][k+2], and kk[k+3][k+3] pairs. (c) Binder cumulant $U_{3}$ for three-fold symmetry and cluster size ratio $d_{\mathrm{cl}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=256$.
Figure 13: Time development of number densities of states in the six-state Potts model at $J_{k,[k+2]}=0$ and $L=128$ (Pot6c). (a) Temporal coexistence of M2W3 and M2HC3 at $h=1.3$ and $J_{k,[k+3]}=2.4$. (b) Temporal coexistence of W3 and HC6 at $h=0.7$ and $J_{k,[k+3]}=-0.6$.
@@ III.1.3 Modes for $J_{k,[k+2]}=0$
Next, we investigate the $J_{k,[k+3]}$ dependence while keeping $J_{k,[k+2]}=0$ (Figs. 1113). We refer to this condition as Pot6c. For $J_{k,[k+3]}<0$, the W3 mode emerges in addition to the HC6, WI, and W6 modes, since the contact between the $k$ and $[k+2]$ states becomes more stable than that between the $k$ and $[k+3]$ states (see Figs. 11(c) and 12). The transition between the W3 and HC6 modes occurs via the temporal coexistence of the two modes, as shown in Fig. 13(b), similar to that occurring for the transition from HC3 to W3 in the three-state Potts model [34, 35]. The ratio of the W3 modes monotonously increases with decreasing $J_{k,[k+3]}$, and we determine the modes from the ratio of the phase periods.
For $J_{k,[k+3]}>0$ two additional modes (M2W3 and M2HC3) are found. At $J_{k,[k+3]}=J_{k,k}=2$, random mixing of two diagonal states ($s=k$ and $[k+3]$) can occur without loss of contact energy in the absence of the other four states. Hence, three types of mixed phases comprising two diagonal states exist with equal stability under $h=0$. For $J_{k,[k+3]}>J_{k,k}$, diagonal-state contacts are preferred over those of the same states, as in the case for antiferromagnetic interactions in spin systems. Thus, for $J_{k,[k+3]}\simeq J_{k,k}$, the homogeneous cycling and wave modes emerge for these three phases at low and high $h$, respectively, like in the three-state Potts model. We call them M2HC3 and M2W3 (i.e., MmmHCnn and MmmWnn for $n$ phases of $m$ mixed phases). The transition between the M2W3 and M2HC3 modes occurs via the temporal coexistence of the two modes (see Fig. 13(a)).
In the M2W3 mode, the three types of mixed domains form spiral waves (see Fig. 11(a) and Movie S4). To characterize this mode, we calculated the $p_{\mathrm{contact}}$ values and cluster size ratio $d_{\mathrm{cl}}$. Clusters of single or two diagonal states are considered. Note that, in the latter case, when neighboring sites have the same state or diagonal state ($s_{i}=k$ and $s_{j}=k$ or $[k+3]$), they belong to the same cluster. With increasing $J_{k,[k+3]}$, the ratio $d_{\mathrm{cl}}=n_{\mathrm{c1}}/n_{\mathrm{c2}}$ of the mean sizes for these two clusters ($n_{\mathrm{c1}}$ and $n_{\mathrm{c2}}$ for the single and two diagonal states, respectively) decreases and $p_{\mathrm{contact}}$ of the diagonal pairs ($s=k$ and $[k+3]$) increases (see Fig. 12(b) and (c)). We consider that the W6 mode changes to the M2W3 mode when the diagonal-pair $p_{\mathrm{contact}}$ exceeds the other probabilities.
Figure 14: Dependence on $J_{k,[k+3]}-J_{k,[k+2]}$ in the six-state Potts model at $J_{k,[k+2]}+J_{k,[k+3]}=1$, $h=1$, and $L=128$. (a) Order parameters $R_{1}$ and $R_{3}$. The bidirectional arrows at the top represent the ranges of dynamic modes, including the coexistence states. (b) Susceptibilities $\chi_{1}$ and $\chi_{3}$. The conditions of $J_{k,[k+3]}-J_{k,[k+2]}=-1$, 0, and 11 correspond to the points in the phase diagram of Figs. 7, 1, and 11 (Pot6b, Pot6a, and Pot6c), respectively. (c) Time development of number densities of states at $J_{k,[k+3]}-J_{k,[k+2]}=-0.4$. The HC3 and W3 modes temporally coexist.
Figure 15: Probabilities $p_{\mathrm{phase}}$ of one-phase dominant and multi-phase coexistence states as functions of $h$ at $J_{k,k^{\prime}}=2\delta_{k,k^{\prime}}$ and $L=128$ for the $q$-state Potts models with (a) $q=8$, (b) $q=5$, and (c) $q=4$. The plot in (c) is reproduced from Ref. 36. Licensed under $C$ BY.
@@ III.2 Five-State Potts Model
When the interactions of the standard Potts model with $q$ states are used ($J_{k,k^{\prime}}=2\delta_{k,k^{\prime}}$), HCqq and Wqq modes are obtained at low and high $h$, respectively, for $3\leq q\leq 8$ (see Fig. 15, the data at $q=7$ are not shown). When $q>4$, the WI mode emerges for an intermediate value of $h$, whereas a direct transition of two modes occurs for $q=3$ and $4$ [34, 36]. Therefore, the dynamic behaviors of the five-state Potts model with $J_{k,[k+2]}=0$ are similar to those of the six-state model with $J_{k,[k+2]}=J_{k,[k+3]}=0$.
The dynamic phase diagram of the five-state Potts model is shown in Fig. 16. Since $5$ is a prime number, no factorized-symmetry modes exist. At $J_{k,[k+2]}\simeq J_{k,k}=2$, a five-state mixed phase (M5) exists, as for the six-state Potts model. To characterize the transition between the M5 and W5 modes, we calculated $p_{\mathrm{phase}}$, $p_{\mathrm{contact}}$, and the mean cluster size $n_{\mathrm{c1}}$, as shown in Fig. 17; At $J_{k,[k+2]}\simeq 0.6$, a stepwise change is observed for $p_{\mathrm{phase}}$ of five states. Additionally, a minimum is observed for $p_{\mathrm{contact}}$ of kk[k+1][k+1] pairs and a sharp decrease in $n_{\mathrm{c1}}$ is initiated. For $J_{k,[k+2]}\lesssim 0.6$, $n_{\mathrm{c1}}$ is dependent on the system size; however, no size dependence is detected for $J_{k,[k+2]}\gtrsim 0.6$. In contrast, no size dependence is obtained for $p_{\mathrm{contact}}$ in either range (compare the curves for $L=128$ and $256$ in Fig. 17(b) and (c)). Thus, we consider that $J_{k,[k+2]}\simeq 0.6$ is the transition point for $h=1$. The minimum of $p_{\mathrm{contact}}$ is used to determine the transition points for other values of $h$, since it is clearer than the other two features.
@@
Figure 17: Dependence on $J_{k,[k+2]}$ in the five-state Potts model at $h=1$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for kk[k+1][k+1] and kk[k+2][k+2] pairs. (c) Mean cluster size $n_{\mathrm{c1}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=256$.